# Relativistic BEC extracted from a complex FRG flow equation

### v1

@@ II FRG flow equation
A remarkable feature in the flow equation (7) is the complexity. For instance, it is readily found that $E_{-}$ becomes imaginary for a large $\mu$. We emphasize, however, that this feature is not specific to the FRG framework. In fact, in relativistic scalar theories, there is a constraint on a mass and chemical potential: $\mu^{2}\leq({\rm mass})^{2}$, where $(\mathrm{mass})$ is the bare mass in the free case and an effective mass under the MFA [12, 26]. This bound comes from the requirement that the thermodynamic potential be real and finite, or that the Bose distribution be positive. A similar requirement should be imposed in the thermodynamics derived from the FRG framework. Therefore, the effective potential $U_{k}(\rho)$, and thus the right hand side of Eq. (7) should be real and finite. For $k>0$, this is respected as long as $E_{-}^{2}>0$. We note that $E_{-}=0$ is not allowed by the finiteness of the right hand side in Eq. (7). This inequality holds if  [^1] The inequality $E_{-}^{2}\geq 0$ is also fulfilled when $\mu^{2}\geq k^{2}+V_{k}^{\prime}+2\rho V_{k}^{\prime\prime}$. Nevertheless, we have numerically confirmed that this condition never holds for our parameter choices as $\bar{m},T,\mu\ll\Lambda$.
@@
Finally, we briefly mention the numerical method to solve Eq. (7). First, it is useful to introduce the dimensionless valuables $t=\log(k/\Lambda)$ and $r=\rho/\Lambda^{2}$. We then employ the fourth-order Runge-Kutta method for the $t$-derivatives from $t=0$ to $t=\log(10^{-4})$ 222 The corresponding value $k=10^{-4}\Lambda$ is much smaller than $k\simeq\bar{m}$, around which our numerical computation of the FRG flow converges well. This choice is only for the sake of drawing Figs. 1 and 2. , and the grid method [36, 24, 37] for the $r$-derivatives. The discretization widths are chosen as $\Delta t\simeq-4.6\times 10^{-7}$ and $\Delta r=10^{-4}$, and the number of the $t$-integration steps is $2.0\times 10^{7}$. Also we have to set a range of $\rho$ for numerical calculations, and the right endpoint $\rho_{\mathrm{right}}$ and the left $\rho_{\mathrm{left}}$. This $\rho_{\mathrm{left}}$ is zero or determined from Eq. (12) or (13), while $\rho_{\mathrm{right}}$ can be arbitrary as far as $\rho_{\mathrm{right}}\gg\bar{m}^{2},\mu^{2},T^{2}$. The derivatives at $\rho=\rho_{\mathrm{left}}$ ($\rho=\rho_{\mathrm{right}}$) endpoint of the $r$-grids are calculated with the forward (backward) differences, but the others with the central difference. In the present analysis such a treatment becomes crucial, since the left endpoint gives the ground state, i.e., $\rho_{0}=\rho_{\mathrm{left}}$; see the argument in the following section. In order to accurately locate $\rho_{0}$, we also interpolate adjacent grids with the cubic-spline.
@@ Appendix A Derivation of the flow equation
Acknowledgements.
K. M. is supported by the Japan Society for the Promotion of Science (JSPS) KAKENHI under Grant No. 24K17052. F. T. acknowledges support from The Moritani Scholarship Foundation and JST SPRING under Grant No. JPMJSP2151.

### v2

@@ II FRG flow equation
A remarkable feature in the flow equation (7) is the complexity. For instance, it is readily found that $E_{-}$ becomes imaginary for a large $\mu$. We emphasize, however, that this feature is not specific to the FRG framework. In fact, in relativistic scalar theories, there is a constraint on a mass and chemical potential: $\mu^{2}\leq({\rm mass})^{2}$, where $(\mathrm{mass})$ is the bare mass in the free case and an effective mass under the MFA [12, 26]. This bound comes from the requirement that the thermodynamic potential be real and finite, or that the Bose distribution be positive. A similar requirement should be imposed in the thermodynamics derived from the FRG framework. Therefore, the effective potential $U_{k}(\rho)$, and thus the right hand side of Eq. (7) should be real and finite. For $k>0$, this is respected as long as $E_{-}^{2}>0$. We note that $E_{-}=0$ is not allowed by the finiteness of the right hand side in Eq. (7). This inequality holds if  [^1] The inequality $E_{-}^{2}>0$ is also fulfilled when $\mu^{2}>k^{2}+V_{k}^{\prime}+2\rho V_{k}^{\prime\prime}$. Nevertheless, we have numerically confirmed that this condition never holds for our parameter choices as $\bar{m},T,\mu\ll\Lambda$.
@@
The necessity of the above conditions is one of the main findings in this paper, unlike the preceding relevant works. For instance, in Refs. [25] and [33], their flow equations are basically the same as ours, but the above conditions are not taken into account. In the former, the imaginary part in the flow equation is dropped by hand. In the latter, the author chooses the parameters of the bare mass and coupling with which the imaginary part disappears. We also emphasize that since $\mu^{2}\leq(\mathrm{mass})^{2}$ is required in order that the thermal Bose distribution is well-defined, the complexity in the FRG flow equation would be specific to the finite-temperature bosonic theory. Indeed, if the momentum phase space were $O(4)$-symmetric unlike ours over $q=(\omega_{n},\boldsymbol{q})$, the flow equation would be always real [18].
@@
Finally, we briefly mention the numerical method to solve Eq. (7). First, it is useful to introduce the dimensionless valuables $t=\log(k/\Lambda)$ and $r=\rho/\Lambda^{2}$. We then employ the fourth-order Runge-Kutta method for the $t$-derivatives from $t=0$ to $t=\log(10^{-4})$ 222 The corresponding value $k=10^{-4}\Lambda$ is much smaller than $k\simeq\bar{m}$, around which our numerical computation of the FRG flow converges well. This choice is only for the sake of drawing Figs. 1 and 2. , and the grid method [36, 24, 37] for the $r$-derivatives. The discretization widths are chosen as $\Delta t\simeq-4.6\times 10^{-7}$ and $\Delta r=10^{-4}$, and the number of the $t$-integration steps is $2.0\times 10^{7}$. In order to accurately locate $\rho_{0}$, we also interpolate adjacent grids with the cubic-spline.
Also we have to set a range of $\rho$ for numerical calculations, and the right endpoint $\rho_{\mathrm{right}}$ and the left endpoint $\rho_{\mathrm{left}}$. This $\rho_{\mathrm{left}}$ is zero or determined from Eq. (12) or (13), while $\rho_{\mathrm{right}}$ can be arbitrary as far as $\rho_{\mathrm{right}}\gg\bar{m}^{2},\mu^{2},T^{2}$. The derivatives at $\rho=\rho_{\mathrm{left}}$ ($\rho=\rho_{\mathrm{right}}$) endpoint of the $r$-grids are calculated with the forward (backward) differences, but the others with the central difference. In the present analysis such a treatment becomes crucial, since the left endpoint gives the ground state, i.e., $\rho_{0}=\rho_{\mathrm{left}}$; see the argument in the following section. We note that for $E_{-}=0$, there emerges a numerical instability because the flow equation (7) diverges. At a finite $k$, we approximately determine $\rho_{\mathrm{left}}$ as the $\rho$ leading to $|E_{-}|/\bar{m}\simeq 10^{-6}$. However, as mentioned above Eq. (13), $E_{-}=0$ for $U_{k\to 0}^{\prime}=0$ yields no numerical instability. Therefore, such a numerical issue of $\rho_{\mathrm{left}}$ never affects the global minimum, which is determined from $U_{k\to 0}^{\prime}=0$.
@@ Appendix A Derivation of the flow equation
Acknowledgments
K. M. is supported by the Japan Society for the Promotion of Science (JSPS) KAKENHI under Grant No. 24K17052. F. T. acknowledges support from the Moritani Scholarship Foundation (No. 22051) and JST SPRING under Grant No. JPMJSP2151.